OpenStudy (anonymous):

Someone Help!! Solve the equation. Check for extraneous solutions. Type your answers in the blanks. Show your work. |4x+3| = 9 + 2x x=_____ or ______

OpenStudy (anonymous):

3

OpenStudy (anonymous):

fan and medal and @ me for more help

OpenStudy (anonymous):

@IQ250

OpenStudy (anonymous):

yo

OpenStudy (anonymous):

I need to understand the steps

OpenStudy (solomonzelman):

\(\large\color{slate}{ |4x+3| = 9 + 2x }\) this equation will give you two possible outcomes: (outcome 1) \(\large\color{slate}{ 4x+3 = 9 + 2x }\) (outcome 2) \(\large\color{slate}{ 4x+3 = -(9 + 2x) }\)

OpenStudy (solomonzelman):

Now, solve each of the 'outcomes'

OpenStudy (solomonzelman):

Lets start here: Do you know what | x | means ?

OpenStudy (solomonzelman):

For example, | 4| is 4 and |-4| is also 4 why? because it is "abolute value" (like distance) which can't be negative

OpenStudy (solomonzelman):

If you have |x|=4 then x can be equal to 4 or -4 to make the statement true

OpenStudy (anonymous):

I do not know what |x| means

OpenStudy (anonymous):

okay so its |-4x-3| ?

OpenStudy (solomonzelman):

And therefore if we have `|4x+3| = 9 + 2x` we are going to split it up like this: 4x+3 = 9 + 2x and 4x+3 = -(9 + 2x) because the absolute value of 4x+3 can be equal to both 9 + 2x and -(9 + 2x)

OpenStudy (solomonzelman):

As you are done reading, rate this to make sense please from 1 (do not get it at all) to 10 (perfectly understood)

OpenStudy (solomonzelman):

(and after you rate it, if we can, I will proceed.)

OpenStudy (solomonzelman):

|x| is a notation for "absolute value of x"

OpenStudy (anonymous):

4

OpenStudy (solomonzelman):

lets re-consider ` | 4 | ` (absolute value of 4) and ` | -4 | ` (absolute value of negative 4) I am sure you have seen a number line before:  and you can certainly think : | 4 | ` (absolute value of 4) as of a distance from 0 to 4 on a number line (which is equal to 4 [units]) | -4 | ` (absolute value of negative 4) as of a distance from 0 to -4 on a number line (which is equal to 4 [units]) (( distance is still 4, just that it is to the other direction)